# Long-time soliton dynamics via a coarse-grained space-time method

### v2

@@ VI.2.2 Schwinger positronium
Figure 16: Dynamics of a Schwinger positronium composed of a stable pair of soliton and antisoliton that oscillate around each other. The pair is initially at a distance $d$ from each other and are moving towards each other, each with speed $u=0.55$. (a) plots the field $\phi_{x}(x,t)$ at early times ($0\leq t\leq 200$) for $g=0.3,\ d=66$. (b) shows $\phi_{x}(x,t)$ at early times for $g=0.3,\ d=22$. (c) shows $\phi_{x}(x,t)$ at early times for $g=0.2,\ d=22$. (d) plots the value of $\phi_{x}(x,t)$ at late times ($3800\leq t\leq 4000$) for $g=0.3,\ d=22$. (e) presents evolution over time of the energies stored inside a finite domain $-16\leq x\leq 16$ with $g=0.3,\ d=22$. (f) zooms into the early evolution of the energies shown in (e), while (g) zooms into the asymptotic times. (h) plots the field $\phi_{x}(x)$ everywhere in space at $t=4000$. (i) also plots $\phi_{x}(x)$ at $t=4000$ with a closed-up view into the range $-400\leq x\leq 400$ around the positronium. (k) shows the closed-up view of the radiation that leaks from the positronium at $t=4000$.
@@
In the framework of 1+1D field theory, positronium can be understood as a stable bound state of a soliton–antisoliton pair oscillating around one another. In the physical 3+1D QED, positronium is not a bound state but rather a two-body resonance with a well-studied finite lifetime [69, 70].
@@
The evolution from two separated solitons to a bound pair is strictly prohibited by the unperturbed SG model. This means that when two solitons in the pure SG model collide, they simply pass through each other. The solitary wave solutions in that theory are solitons according to the stricter definition of the term  [^4] In the discussion so far we have not made a distinction between a soliton solution and a solitary wave solution. The former is a narrower class of solutions than the latter, that require stricter asymptotic conditions in time. For more information see Ref. [81]. The inclusion of the dynamical mass term $g^{2}\varphi$, however, breaks the integrability of the SG model and may facilitate the formation of a bound positronium from two initially separated solitons.
@@
In Figs. 16a-c we present the early-time space-time dynamics of the charge density $\phi_{x}(x,t)$ that corresponds to a soliton-antisoliton pair for different values for $g$ and the initial separation $d$. The initial condition for $\phi$ in all three cases have been chosen as the analytically known fluxon-antifluxon solution in Eq.20. Comparing Figs. 16a and 16b, one can see that smaller $d$ leads to faster decay in the oscillation of the pair. The reason is clear – a smaller initial separation reduces the initial energy stored in the electric field, which is confined to the space between the two solitons, thereby limiting the energy available to sustain the oscillations that emit radiation. Comparing Figs. 16b with Figs. 16c, one sees that smaller $g$ also results in slower decay. This occurs because the soliton accelerates when propagating in a large-$g$ background, as we also observed in the case of the atom with a fixed charge. The increased acceleration leads to enhanced radiation emission from the soliton, causing the positronium to lose energy more rapidly.
@@
In all of the three cases presented in Figs. 16a-c, although the oscillations decay over time, the pair is not completely annihilated but is reduced to a steady state of persistent oscillation. We show an example in Fig. 16d, where the late-time dynamics for $g=0.3,\ d=22$ (the same parameters as Figs. 16b) is plotted. The oscillations of the pair are stabilized as the decay rate becomes progressively small. The binding of an initially separated SG soliton-antisoliton pair into a breather may remind us of the vortex-antivortex interaction in a lossy junction that was discussed in Sec V.2.2. However, unlike in the lossy junction case, where the breather decays until it vanishes entirely due to the effect of the resistive term $\partial_{t}\varphi$, here the solitons only decay until the system reaches a steady state of self-sustained oscillation. In this massive Schwinger process, the breather seen in Fig. 16d is an oscillating electric dipole, and the internal oscillation of the breather at asymptotic time is facilitated by the attractive force between its two oppositely charged halves.
@@
The evolution of energy within the finite domain $-16\leq x\leq 16$, shown in Fig. 16e, offers further insight into these observations. It can be seen that, following an initial rapid decay, the total energy within this region stabilizes at a steady-state value at late times, indicating that energy leakage from the atom eventually ceases. As shown in Fig. 16g, which provides a close-up view of the energy dynamics in the steady state, the kinetic energy oscillates $\pi$ out of phase with the other energy components. This phase relationship suggests a coherent exchange of energy between the kinetic term, the potential energy of the positronium, and the energy stored in the electric field.
@@
Figure 17: Possible dynamics of a system with large effective mass ($g=0.4$) and is initially composed of a soliton-antisoliton pair moving towards each other at the speed $u=0.55$. The early dynamics of the field $\phi_{x}$ are shown at (a), (b), (c), and (d) for initial separations of the pair $d=22,\ d=16.5,\ d=13.75,$and $d=11$, respectively. Late-time dynamics of $\phi_{x}$ are shown at (e) and (f) for $d=22$ and $d=16.5$, respectively.
@@
Figure 18: Fractalization of the initial velocity into intervals that correspond to either scattered solitons or a bound positronium in the long-time limit. Here, $g\!=\!0.32,$ and $d/u\!=\!20$.
@@ VI.2.3 Long-time phase diagram of 1+1D Positronium
Our numerical experiments indicate that for sufficiently large $g$, the fate of an electron–positron pair — whether they scatter or form a stable bound state — depends sensitively on the initial conditions. In Fig. 18, we map the final state as a function of the initial velocity $u$ ($0.05\leq u\leq 0.7$). Interestingly, there is no single critical velocity separating the scattering and bound states; instead, we find a series of disjoint intervals where the system alternates between these outcomes. This “fractal” dependence on the incident velocity closely resembles the complex behavior seen in $Z_{2}$ soliton dynamics within $\varphi^{4}$ theory [72, 73].
By comparing the dynamics in Fig. 17c where $d\!=\!13.75$ with those in Fig. 17b where $d\!=\!16.5$, we can predict a phase transition at $d$ between these two values where the transition is from a system that sustains a localized atom at steady state to one that does not. Along the same line of thought, by comparing the dynamics in Fig. 17a (where $g\!=\!0.4$) with those in Fig. 16b (where $g\!=\!0.3$) - two dynamics that have the exact same initial conditions for the SG soliton pair but with different values for $g$ we can predict a phase transition at $g\!=\!g_{c}$ somewhere between these two values that separates a phase ($g\!<\!g_{c}$) in which the initial SG soliton pair turns into a stable positronium with a phase ($g\!>\!g_{c}$) in which the initial solitons turn into breathers. To quantitatively identify in parameter space $(u,d,g)$ the boundaries of the possible phases that an initial pair of solitons can evolve into at steady state, further detailed analysis will need to be carried out in the future.
@@ VII Discussion
Secondly, our results motivate further investigation into the analog simulation of relativistic 1+1D QED – the massive Schwinger model – using Josephson Transmission Lines (JTLs). While the present study operates strictly within the confines of the semiclassical regime of the quantum Sine-Gordon model, earlier foundational work [58, 59], has established that classical field dynamics can serve as a meaningful starting point for probing quantum phenomena beyond the reach of perturbative techniques, such as those we investigate in this work. In this context, it is essential for future work to clarify which aspects of the massive Schwinger model can be faithfully captured through analog implementations in JTL architectures.
Thirdly, extending the framework developed here to the associated quantum field theory holds the potential to yield novel insights into non-perturbative phenomena that underpin the foundations of quantum theory — namely, the stability and structure of atoms. Atomic spectra and radiative transition rates are traditionally understood through perturbative expansions in the gauge-field coupling and the renormalization framework [74]. Even the stability of the Hydrogen atom’s ground state has only been rigorously established within perturbative QED through the ingenious use of inequalities [75]. A non-perturbative numerical approach capable of accessing such properties – even in the reduced dimensional setting of 1+1D – would represent a significant advance, offering a unique window into atomic structure that remains out of reach for perturbative approaches to quantum gauge theory.
@@ Appendix A Derivation of the SG equation from the gauge-invariant EHDS equations
In this appendix, we show that the sine-Gordon equation can be derived from an electrohydrodynamics description of superconductors (EHDS) [43, 76] applied to the long Josephson junction. This will confirm that although the SG equation is a 1D effective field equation in which the spatial variation along the thickness of the junction is not considered, it can still be rigorously traced back to a first principles formulation.
Consider a long Josephson junction composed of two superconducting islands that are separated by a thin ($\lesssim$ few $nm$) layer of either insulator or normal metal, as illustrated in Fig. 19. The long dimension of the junction is along $x$ and has length $L\!\gg\!\lambda_{J}$, where $\lambda_{J}$ is the Josephson penetration depth. The superconductor-insulator-superconductor variation in material properties is along $z$, while the remaining dimension $W$ of the junction along $y$ is assumed to be much less than $\lambda_{J}$ so that there is effectively no field variation inside the junction in this direction. If the separation between the superconducting island is thin, the superconducting condensate can leak into this region and, if the decay tails from two islands touch, a supercurrent can be established in the junction [77]. Let $\psi$ be the macroscopic order parameter of the collective superconducting condensate fluid that lives in the entire junction and is made of Cooper pairs (with mass $m\!=\!2m_{e}$ and charge $q\!=\!2e$, where $m_{e}$ and $e$ are the electron mass and charge respectively). Using the Madelung representation $\psi=\sqrt{\rho}e^{i\theta}$, where $\rho$ is the local density of condensate and $\theta$ is the gauge-dependent superconducting phase, the EHDS equation governing the dynamics of the condensate in the presence of the electromagnetic field $(V,{\bf A})$ is given by
××${\bf\nabla}\!\times\!{\bf\nabla}\!\times\!\bm{\mathcal{A}}+\mu_{0}$ $\epsilon\frac{\partial^{2}\bm{\mathcal{A}}}{\partial t^{2}}+\frac{\mu_{0}q^{2}}{m}\rho\bm{\mathcal{A}}-\frac{\mu_{0}\epsilon q}{2m}\frac{\partial}{\partial t}\nabla\big{|}\bm{\mathcal{A}}\big{|}^{2}$
@@
where $\bm{\mathcal{A}}={\bf A}-\frac{\hbar}{q}\nabla\theta$ is the gauge-invariant hybridized field that lives everywhere throughout space. When applied to a short Josephson junction with a sufficiently thin separation between the superconducting islands, it was rigorously shown in Ref. [43] that the Josephson current-phase relation is produced from the last two nonlinear terms on the left hand side of Eq. A canceling each other. For a long Josephson junction, the same cancellation holds if the spatial variations of the field $\bm{\mathcal{A}}$ along $x$ and $y$ appear at much larger length scales than along $z$. In the case of Josephson fluxons, the fluxons are assumed to be translationally invariant throughout the $y$ direction. The sizes of the fluxons along $x$ are determined by the Swihart velocity and the speed of the fluxons. These conditions typically result in fluxon sizes on the order of $10\!-\!100$ nm. Therefore, the condition for the mentioned cancellation of two nonlinear terms is well satisfied.
@@
where we have used the sinusoidal dependence on $\varphi$ to express the supercurrent. By rewriting the first term of Eq. A above into $-\partial_{x}^{2}\varphi$, we obtain the sine-Gordon equation.
@@ Appendix B Bosonization of the massive Schwinger model
More precisely, employing the partition function analysis by Haldane [78], one can prove the following operator identity:
@@
$\begin{split}\phi(t,x)\equiv\sum_{\eta}\phi^{\eta}(t,x)=\frac{e}{\sqrt{\pi}}E(t,x)\end{split}$ (37)
In the $L\rightarrow\infty$ limit, Eq. 36 and Eq. 37 allow us to rewrite Eq. 34 as the following Lagrangian density in terms of $\phi(t,x)$ alone:
$\begin{split}\mathcal{L}=&\frac{1}{2}\partial_{\mu}\phi(t,x)\partial^{\mu}\phi(t,x)-\frac{e^{2}}{2\pi}\phi^{2}(t,x)\\ &+\frac{c}{\pi\epsilon}\cos\big(2\sqrt{\pi}\phi(t,x)\big{missing})\end{split}$ (38)
@@
$m\bar{\psi}\psi\to-\frac{c}{\pi\epsilon}\cos(2\sqrt{\pi}\phi)$ (39)
@@
$\mathcal{L}=\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\varphi+\frac{\kappa}{2}\cos(2\sqrt{\pi}\varphi)+\frac{g}{2}\varphi\epsilon^{\mu\nu}F_{\mu\nu}-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ (40)
with $\kappa=\frac{e^{\gamma}me}{\pi^{3/2}}$ for small fermion masses. In addition, one obtains a $\phi^{2}$ term in the bosonic Lagrangian due to the gauge interaction, which gives rise to a free Schwinger boson mass of $g=\frac{e}{\sqrt{\pi}}$.

### v3

@@ VI.2.2 Schwinger positronium
Figure 16: Dynamics of a Schwinger positronium composed of a stable pair of soliton and antisoliton that oscillate around each other. The pair is initially at a distance $d$ from each other and are moving towards each other, each with speed $u=0.55$. (a) plots the field $\phi_{x}(x,t)$ at early times ($0\leq t\leq 200$) for g=0.3,d=66g=0.3,\ d=66. (b) shows $\phi_{x}(x,t)$ at early times for g=0.3,d=22g=0.3,\ d=22. (c) shows $\phi_{x}(x,t)$ at early times for g=0.2,d=22g=0.2,\ d=22. (d) plots the value of $\phi_{x}(x,t)$ at late times ($3800\leq t\leq 4000$) for g=0.3,d=22g=0.3,\ d=22. (e) presents evolution over time of the energies stored inside a finite domain $16\leq x\leq 16$ with g=0.3,d=22g=0.3,\ d=22. (f) zooms into the early evolution of the energies shown in (e), while (g) zooms into the asymptotic times. (h) plots the field $\phi_{x}(x)$ everywhere in space at $t=4000$. (i) also plots $\phi_{x}(x)$ at $t=4000$ with a closed-up view into the range $400\leq x\leq 400$ around the positronium. (k) shows the closed-up view of the radiation that leaks from the positronium at $t=4000$.
@@
In the framework of 1+1D field theory, positronium can be understood as a stable bound state of a soliton–antisoliton pair repeatedly passing through each other as they oscillate around a point $x$. In the physical 3+1D QED, positronium is not a bound state but rather a two-body resonance with a well-studied finite lifetime [karshenboim2004precision, badertscher2007].
@@
The evolution from two separated solitons to a bound pair is strictly prohibited by the unperturbed SG model. This means that when two solitons in the pure SG model collide, they simply pass through each other. The solitary wave solutions in that theory are solitons according to the stricter definition of the term  [^4] In the discussion so far we have not made a distinction between a soliton solution and a solitary wave solution. The former is a narrower class of solutions than the latter, that require stricter asymptotic conditions in time. For more information see Ref. [ablowitz1981solitons]. The inclusion of the dynamical mass term $g^{2}\varphi$, however, breaks the integrability of the SG model and may facilitate the formation of a bound positronium from two initially separated solitons.
@@
In Figs. 16a-c we present the early-time space-time dynamics of the charge density $\phi_{x}(x,t)$ that corresponds to a soliton-antisoliton pair for different values for $g$ and the initial separation $d$. The initial condition for $\varphi$ in all three cases have been chosen as the analytically known fluxon-antifluxon solution in Eq.21. Comparing Figs. 16a and 16b, one can see that smaller $d$ leads to faster decay in the oscillation of the pair. The reason is clear – a smaller initial separation reduces the initial energy stored in the electric field, which is confined to the space between the two solitons, thereby limiting the energy available to sustain the oscillations that emit radiation. Comparing Figs. 16b with Figs. 16c, one sees that smaller $g$ also results in slower decay. This occurs because the soliton accelerates when propagating in a large-$g$ background, as we also observed in the case of the atom with a fixed charge. The increased acceleration leads to enhanced radiation emission from the soliton, causing the positronium to lose energy more rapidly.
@@
In all of the three cases presented in Figs. 16a-c, although the oscillations decay over time, the pair is not completely annihilated but is reduced to a steady state of persistent oscillation. We show an example in Fig. 16d, where the late-time dynamics for g=0.3,d=22g=0.3,\ d=22 (the same parameters as Figs. 16b) is plotted. The oscillations of the pair are stabilized as the decay rate becomes progressively small. The binding of an initially separated SG soliton-antisoliton pair into a breather may remind us of the vortex-antivortex interaction in a lossy junction that was discussed in Sec V.2.2. However, unlike in the lossy junction case, where the breather decays until it vanishes entirely due to the effect of the resistive term $\partial_{t}\varphi$, here the solitons only decay until the system reaches a steady state of self-sustained oscillation. In this massive Schwinger process, the breather seen in Fig. 16d is an oscillating electric dipole, and the internal oscillation of the breather at asymptotic time is facilitated by the attractive force between its two oppositely charged halves.
@@
The evolution of energy within the finite domain $16\leq x\leq 16$, shown in Fig. 16e, offers further insight into these observations. It can be seen that, following an initial rapid decay, the total energy within this region stabilizes at a steady-state value at late times, indicating that energy leakage from the atom eventually ceases. As shown in Fig. 16g, which provides a close-up view of the energy dynamics in the steady state, the kinetic energy oscillates $\pi$ out of phase with the other energy components. This phase relationship suggests a coherent exchange of energy between the kinetic term, the potential energy of the positronium, and the energy stored in the electric field.
@@
Figure 17: Possible dynamics of a system with large effective mass ($g=0.4$) and is initially composed of a soliton-antisoliton pair moving towards each other at the speed $u=0.55$. The early dynamics of the field $\phi_{x}$ are shown at (a), (b), (c), and (d) for initial separations of the pair d=22,d=16.5,d=13.75,d=22,\ d=16.5,\ d=13.75,\ and $d=11$, respectively. Late-time dynamics of $\phi_{x}$ are shown at (e) and (f) for $d=22$ and $d=16.5$, respectively.
@@
Figure 18: Fractalization of the initial velocity into intervals that correspond to either scattered solitons or a bound positronium in the long-time limit. Here, g=0.32,g\!=\!0.32, and d/u=20d/u\!=\!20.
@@ VI.2.3 Long-time phase diagram of 1+1D Positronium
Our numerical experiments indicate that for sufficiently large $g$, the fate of an electron–positron pair — whether they scatter or form a stable bound state — depends sensitively on the initial conditions. In Fig. 18, we map the final state as a function of the initial velocity $u$ ($0.05\leq u\leq 0.7$). Interestingly, there is no single critical velocity separating the scattering and bound states; instead, we find a series of disjoint intervals where the system alternates between these outcomes. This “fractal” dependence on the incident velocity closely resembles the complex behavior seen in $Z_{2}$ soliton dynamics within $\varphi^{4}$ theory [anninos1991fractal, campbell1983resonance].
By comparing the dynamics in Fig. 17c where d=13.75d\!=\!13.75 with those in Fig. 17b where d=16.5d\!=\!16.5, we can predict a phase transition at $d$ between these two values where the transition is from a system that sustains a localized atom at steady state to one that does not. Along the same line of thought, by comparing the dynamics in Fig. 17a (where g=0.4g\!=\!0.4) with those in Fig. 16b (where g=0.3g\!=\!0.3) - two dynamics that have the exact same initial conditions for the SG soliton pair but with different values for $g$ - we can predict a phase transition at $g\!=\!g_{c}$ somewhere between these two values that separates a phase ($g\!<\!g_{c}$) in which the initial SG soliton pair turns into a stable positronium with a phase ($g\!>\!g_{c}$) in which the initial solitons turn into breathers. To quantitatively identify in parameter space $(u,d,g)$ the boundaries of the possible phases that an initial pair of solitons can evolve into at steady state, further detailed analysis will need to be carried out in the future.
@@ VII Discussion
Secondly, our results motivate further investigation into the analog simulation of relativistic 1+1D QED – the massive Schwinger model – using Josephson Transmission Lines (JTLs). While the present study operates strictly within the confines of the semiclassical regime of the quantum Sine-Gordon model, earlier foundational work [dashen1974, dashen1975], has established that classical field dynamics can serve as a meaningful starting point for probing quantum phenomena beyond the reach of perturbative techniques, such as those we investigate in this work. In this context, it is essential for future work to clarify which aspects of the massive Schwinger model can be faithfully captured through analog implementations in JTL architectures.
Thirdly, extending the framework developed here to the associated quantum field theory holds the potential to yield novel insights into non-perturbative phenomena that underpin the foundations of quantum theory — namely, the stability and structure of atoms. Atomic spectra and radiative transition rates are traditionally understood through perturbative expansions in the gauge-field coupling and the renormalization framework [weinberg2005qft]. Even the stability of the Hydrogen atom’s ground state has only been rigorously established within perturbative QED through the ingenious use of inequalities [lieb2005stability]. A non-perturbative numerical approach capable of accessing such properties – even in the reduced dimensional setting of 1+1D – would represent a significant advance, offering a unique window into atomic structure that remains out of reach for perturbative approaches to quantum gauge theory.
@@ Appendix A Derivation of the SG equation from the gauge-invariant EHDS equations
In this appendix, we show that the sine-Gordon equation can be derived from an electrohydrodynamics description of superconductors (EHDS) [dec-qed, maldonado2024pressure] applied to the long Josephson junction. This will confirm that although the SG equation is a 1D effective field equation in which the spatial variation along the thickness of the junction is not considered, it can still be rigorously traced back to a first principles formulation.
Consider a long Josephson junction composed of two superconducting islands that are separated by a thin ($\lesssim$ few $nm$) layer of either insulator or normal metal, as illustrated in Fig. 19. The long dimension of the junction is along $x$ and has length $L\!\gg\!\lambda_{J}$, where $\lambda_{J}$ is the Josephson penetration depth. The superconductor-insulator-superconductor variation in material properties is along $z$, while the remaining dimension $W$ of the junction along $y$ is assumed to be much less than $\lambda_{J}$ so that there is effectively no field variation inside the junction in this direction. If the separation between the superconducting island is thin, the superconducting condensate can leak into this region and, if the decay tails from two islands touch, a supercurrent can be established in the junction [JosephsonReview_1965]. Let $\psi$ be the macroscopic order parameter of the collective superconducting condensate fluid that lives in the entire junction and is made of Cooper pairs (with mass $m\!=\!2m_{e}$ and charge q=2eq\!=\!2e, where $m_{e}$ and $e$ are the electron mass and charge respectively). Using the Madelung representation $\psi=\sqrt{\rho}e^{i\theta}$, where $\rho$ is the local density of condensate and $\theta$ is the gauge-dependent superconducting phase, the EHDS equation governing the dynamics of the condensate in the presence of the electromagnetic field $(V,{\bf A})$ is given by
${\bf\nabla}\!\times\!{\bf\nabla}\!\times\!\bm{\mathcal{A}}+\mu_{0}$ $\epsilon\frac{\partial^{2}\bm{\mathcal{A}}}{\partial t^{2}}+\frac{\mu_{0}q^{2}}{m}\rho\bm{\mathcal{A}}-\frac{\mu_{0}\epsilon q}{2m}\frac{\partial}{\partial t}\nabla\big|\bm{\mathcal{A}}\big|^{2}$
@@
[…]
@@
where we have used the sinusoidal dependence on $\varphi$ to express the supercurrent. By rewriting the first term of Eq. A above into $\partial_{x}^{2}\varphi$, we obtain the sine-Gordon equation.
@@ Appendix B Bosonization of the massive Schwinger model
More precisely, employing the partition function analysis by Haldane [Haldane_1981], one can prove the following operator identity:
@@
$\begin{split}\varphi(t,x)\equiv\sum_{\eta}\phi^{\eta}(t,x)=\frac{e}{\sqrt{\pi}}E(t,x)\end{split}$ (38)
In the $L\rightarrow\infty$ limit, Eq. 37 and Eq. 38 allow us to rewrite Eq. 35 as the following Lagrangian density in terms of $\varphi(t,x)$ alone:
[…]
@@
$m\bar{\psi}\psi\to-\frac{c}{\pi\epsilon}\cos(2\sqrt{\pi}\varphi)$ (40)
@@
One obtains a $\varphi^{2}$ term in the bosonic Lagrangian due to the gauge interaction, which gives rise to a free Schwinger boson mass of $g$. The new coefficients $g$, $\kappa$ are obtained after applying the bosonization dictionary and normal ordering with respect to the renormalized boson modes. They are defined as
$g\equiv\frac{e}{\sqrt{\pi}}$ (42)
$\kappa\equiv\frac{mee^{\gamma}}{\pi^{3/2}}\Bigg(\frac{e^{\gamma}m}{e/\sqrt{\pi}}+\sqrt{1+\frac{(e^{\gamma}m)^{2}}{e^{2}/\pi}}\Bigg)$ (43)
In the limit of small $\frac{m}{e}$, we obtain the usual expression for $\kappa\equiv\frac{e^{\gamma}me}{\pi^{3/2}}$.